# What Lurks Within? Concept Auditing for Shared Diffusion Models at Scale

### v1

@@ 2.2. Parameter-Efficient Fine-Tuning (PEFT)
3. Problem Setup: Concept Auditing in the Wild
While the widespread adoption of PEFT has enabled individual users to rapidly customize and share DMs across public platforms, it also introduces serious risks: many fine-tuned models are trained on unverified, proprietary, or sensitive datasets, with minimal oversight or transparency. This raises an urgent need for concept auditing: verifying whether a shared DM has learned to generate specific high-risk concepts.
@@ 3. Problem Setup: Concept Auditing in the Wild
These challenges render the current auditing pipelines, whose reliance on observable behaviors—prompts and outputs, brittle, expensive, and unreliable in real-world auditing scenarios. Therefore, we ask a more direct and scalable question: what the model has internally learned, regardless of how it is prompted or what it outputs? This motivates the need for a model-centric auditing framework that avoids reliance on prompts or outputs and instead operates directly on the model’s internal behavior. We present our solution in the next section.
4. PAIA: A Model-Centric Auditing Framework
@@ 4.1.2. Theoretical Insight
We provide a full proof in Appendix (Section A).
@@
Figure 2. Impact of prompts on cross-attention. We present the results of attention map $\|S\|$ and $\|diag(S)-SS^{T}\|$ with respect to the first token [BOS] in the prompt. The results are averaged over all models in two categories in our experiments, detailed in Section 5.1.
@@
The gradient of the cross-attention function diminishes diminishes as the generation progresses (i.e., as $t$ becomes smaller).
@@ 4.1.3. Empirical Analysis
To complement this theoretical finding, we empirically examine how prompt influence evolves during the denoising process. Specifically, we track two metrics across timesteps: (1) the magnitude of the attention map $|S|$, and (2) the value of $|diag(S)-SS^{T}|$, which reflects prompt sensitivity. We focus on the [BOS] token, which typically carries the highest semantic weight in a prompt. As shown in Figure 2, during the later stages of generation (smaller $t$), the attention map becomes more focused, with $|S|$ approaching $1$. Simultaneously, $|diag(S)-SS^{T}|$ decreases toward 00, indicating a lower sensitivity to prompt variations. This empirical trend confirms our theoretical result: in the later stages of the generation process, the model becomes increasingly confident about the image contents and relies less on the conditioning text prompt.
@@ 4.2.1. Calibrated Error Measurement
where $\mathcal{D}_{target}$ represents the images of the target concept.
Figure 3. Calibration error across denoising timesteps. We compare the difference in denoising error between the fine-tuned model and its base model for target vs. irrelevant concepts. Early stages (large $t$) offer stronger separation but are more prompt-sensitive. This motivates our conditional calibrated error (CCE), which preserves early-stage signals while reducing prompt influence.
@@
If the calibrated error falls below the threshold, the detector predicts that the fine-tuned DM can generate the target concept. Figure 3 demonstrates the effectiveness of this approach: the calibrated errors for target concepts are significantly lower than those for irrelevant concepts, enabling accurate differentiation. Note that to ensure the analysis is not influenced by inaccurate prompts, we employ the original (accurate) prompts used during fine-tuning. The calibration error is normalized by the denoising error of the base model to account for variations across different time steps, ensuring a fair comparison.
@@ 4.2.2. Conditional Calibrated Error Measurement
While calibrated error effectively identifies target concepts, it exhibits varying performance across different stages of the DM generation process. As shown in Figure 3, the calibration errors (using original prompts) in the early stages of generation (large $t$) are more effective in concept identification. This finding aligns with the recent studies on DM generation mechanisms (zhang2024cross, ; yi2024towards, ): DMs primarily generate semantic information (e.g., structure) during early stages, while later stages focus on refining image details (e.g., texture). Since semantic information is more directly tied to concepts, later-stage details alone may be insufficient for distinguishing between concepts.
@@ 4.4. Overall PAIA Pipeline
We present the overall framework of PAIA, which integrates the prompt-agnostic and image-free design principles, along with unsupervised concept detector. The complete procedure of PAIA consists of the following four steps.
  • Step 1: Construct Irrelevant Concept Set. Collect a set of images representing concepts the DM is not expected to generate. For each image, compute conditional calibrated error (CCE) $\mathcal{L}_{cce}^{t}$ (Eq. 14) across multiple timesteps. CCE integrates early-stage measurements with frozen cross-attention parameters and late-stage measurements using all fine-tuned parameters.
@@
  • Step 4: Detect Concept Generation. Apply the trained outlier detector to the target concept’s calibrated errors. If these values are statistically distinct from the baseline (i.e., flagged as outliers), the model concludes that the DM has learned to generate the target concept.
@@ 5.1. Experimental Settings
We first evaluate the performance of PAIA on two categories of fine-tuned DMs and compare it against six baselines. Details on dataset construction, image examples, and random prompt term lists are provided in Appendix D.
@@ 5.1.1. Datasets and Fine-tuned DMs
We generate textual prompts for each image using BLIP (li2022blip, ). For auditing, we define two trigger types per concept: a normal trigger that aligns with the target concept, and an abnormal trigger that is semantically unrelated, simulating the effect of a simple backdoor-style trigger. Datasets are split evenly into training and testing subsets; training data is used for LoRA fine-tuning, and test data is used for auditing evaluation. Full sample visualizations are included in Appendix D.
All DMs are fine-tuned from Stable Diffusion 1.5555https://huggingface.co/stabilityai/stable-diffusion-1-5 using Low-Rank Adaptation (LoRA), with rank $r=64$, a learning rate of $1e$-$4$, and the AdamW optimizer. Each model is trained for $50$ epochs. For each category, we fine-tune models with $1$, $2$, and $3$ concepts per model to evaluate multi-concept auditing. This results in $50$ celebrity and $10$ cartoon models for each setting. Additionally, we include $50$ celebrity and $10$ cartoon models trained with abnormal (i.e., semantically unrelated) trigger words to evaluate robustness under mismatched prompt conditions.
@@ 5.1.2. PAIA Settings
Outlier detection model. In PAIA, we deploy a widely used data anomaly detection algorithm, Isolation Forest (liu2008isolation, ), as the detector. We train the Isolation Forest model on the irrelevant images (the DM model cannot generate them) and then predict if the images of the target concept are irrelevant. If not, then we predict the DM can generate this concept and vice versa.
The outlier detection model is trained with $100$ images with irrelevant concepts. In our evaluation, we randomly select images from the same category (celebrity or cartoon) but with different concepts. In practice, a straightforward strategy for constructing such a set is to use rare images —those that are unlikely to be generated by fine-tuned DMs.
@@
  • Random. The prompt is composed of five randomly selected terms from a common word list for generating DMs, e.g., “natural lighting,” “best quality,” “ultra detailed.” A complete list is presented in Table 8 in Appendix.
  • Null. The null text, i.e., “”, is employed as the prompt input for DMs.
@@ 5.2. Baselines
  • Naive: We consider a naive probing approach that uses an image captioning model, GIT (wang2022git, ), to derive prompts from the target images.
@@
  • Image Classifier: We train a ConvNeXt (liu2022convnet, )-based image classifier to identify celebrity/cartoon concepts. The classifier is pre-trained on ImageNet and fine-tuned on the collected training data. After fine-tuning, the image classifier achieves $89.2$% and $98.1$% accuracy on celebrity and cartoon data, respectively.
@@ 5.3. Comparison with Baselines
We compare the auditing performance between PAIA and baselines. We first show the results on normal triggers, where the triggers relevant to the target concept are used in fine-tuning DMs, and then the results on abnormal triggers, which are irrelevant to the target concept.
Table 1. Performance comparison on Celebrity DMs. We report accuracy, precision, recall, F1 score, and auditing time for detecting a concept with $10$ images.
@@
Performance Comparison on Normal Triggers: We first compare PAIA with baselines with results presented in Tables 1 and 2. PAIA consistently achieves the highest accuracy (92% and 91%), precision (90% and 89%), and recall (94% for both categories), demonstrating its superior performance in concept auditing. Among the baselines, the PEZ Classifier performs relatively well with accuracy of 75% and 77%, but still significantly lower than PAIA. Other methods, such as MU Classifier and MU CLIP, achieve high recall but suffer from poor precision and accuracy, indicating a high positive rate.
@@
Compared to baselines, PAIA demonstrates much greater efficiency and scalability. Prompt probing techniques (MU and PEZ) often require considerably longer optimization times. Even against naive image-captioning-based methods, PAIA is more efficient in the second scenario since it computes denoising errors at a limited number of time steps rather than performing the full generation process.
Performance Comparison on Abnormal Triggers: The results for abnormal triggers follow a similar trend to those observed with normal trigger words. PAIA consistently outperforms all baseline methods across both datasets, achieving the highest accuracy, precision, recall, and F1 score. PEZ Classifier shows competitive performance across baselines, while methods like MU Classifier and MU CLIP continue to suffer from high false positive rates, leading to poor precision and low overall accuracy. Interestingly, these methods perform even worse than the naive prompt probing approaches, such as Naive Classifier and Naive CLIP, which do not rely on sophisticated optimization. This suggests that the MU methods face significant challenges when attempting to optimize for abnormal triggers. The results underscore the robustness of PAIA, which remains effective in detecting concepts fine-tuned on both normal and abnormal triggers.
An interesting observation is that auditing performance on abnormal triggers is generally better than on normal triggers. We attribute this to differences in the fine-tuned DMs rather than the detectors themselves, likely due to stronger memorization of abnormal triggers in fine-tuned DMs, a phenomenon noted in backdoor attack research. For instance, Naive Classifier and Naive CLIP demonstrate improved auditing performance on abnormal triggers, despite not utilizing any prompt information. This suggests that the difference arises from the performance of the fine-tuned DMs rather than the concept auditing methods.
Table 3. Performance comparison on Celebrity DMs with abnormal triggers.
Detector Accuracy Precision Recall F1 Score
Naive Classifier 75% 86% 60% 71%
Naive CLIP 77% 81% 70% 75%
MU Classifier 57% 54% 100% 70%
MU CLIP 57% 54% 100% 70%
PEZ Classifier 82% 81% 84% 82%
PEZ CLIP 65% 60% 92% 72%
PAIA 94% 89% 100% 94%
Table 4. Performance comparison on Cartoon DMs with abnormal triggers.
Detector Accuracy Precision Recall F1 Score
Naive Classifier 85% 86% 84% 85%
Naive CLIP 55% 55% 58% 56%
MU Classifier 56% 53% 100% 69%
MU CLIP 45% 47% 86% 61%
PEZ Classifier 83% 75% 100% 85%
PEZ CLIP 53% 52% 72% 61%
PAIA 92% 86% 100% 93%
@@ 5.4. Ablation Study
Effectiveness of Conditional Calibrated Error Measurement: We analyze the role of conditional calibrated error measurement in Figures 5 and 6. We compare the auditing performance across four settings: 1) freezing cross-attention layers at the early stages (i.e., conditional calibrated error), 2) not freezing any layers (i.e., calibrated error), 3) freezing cross-attention layers at all stages, and 4) freezing self-attention layers at all stages. The results show that by freezing parameters in cross-attention layers at the early stages of generation, conditional calibration error improves the performance of PAIA in most cases.
@@
Effectiveness of Outlier Detection Algorithms: We investigate the effectiveness of different outlier detection methods, including Isolation Forest (IF) (liu2008isolation, ), Angle-based outlier detection (ABOD) (kriegel2008angle, ), k-Nearest Neighbors (kNN) (laaksonen1996classification, ), Gaussian Mixture Model (GMM) (aggarwal2017introduction, ), and One-class SVM (OneSVM) (manevitz2001one, ). We evaluate these algorithms on both celebrity and cartoon DMs with 1 concept and normal triggers. Among the algorithms investigated, Isolated Forest (the default algorithm in PAIA) consistently achieved the highest accuracy and F1 scores, but other robust outlier detection algorithms like ABOD and kNN also performed competitively, with slightly lower precision.
@@
Impact of Number of Irrelevant Images: PAIA computes the conditional calibration errors from irrelevant images and utilizes these errors to train the outlier detector. We examine how the number of irrelevant images impacts auditing performance, as shown in Figure 9. While irrelevant images (from concepts that the fine-tuned DMs cannot generate) are generally accessible to the detector, increasing their number can lead to a longer auditing time. By default, we use $200$ irrelevant images. The results indicate that using $120$ images is sufficient for training an effective detector in PAIA. Adding more irrelevant images beyond this threshold does not significantly improve auditing performance.
Figure 9. Impact of the number of irrelevant images for training the detector.

### v2

@@ 2.2. Parameter-Efficient Fine-Tuning (PEFT)
3. PAIA: A Model-Centric Auditing Framework
3.1. Threat Model
We assume the auditor can access parameters from both the fine-tuned and the base model, but not know how the model is fine-tuned, i.e., what concepts are involved in the fine-tuning and what training data and prompts are used. This assumption is practical and common for model-sharing platforms (e.g., Civitai, HuggingFace). These platforms store model parameters for distribution and, in the event of DMCA takedown requests (e.g., for copyright infringement), can access both the fine-tuned and base models for analysis. Note that we don’t assume a typical black-box setting (e.g., input–output access only), as it is not representative in concept auditing scenarios. In addition, we assume the auditor has access to the sample images of the target concept, which will be used to evaluate whether the DMs can generate such concept. This assumption is practical, e.g., in a takedown request for a cartoon character, the requester would provide sample images of the target character.
3.2. Motivation and Challenges in the Wild
While the widespread adoption of PEFT has enabled individual users to rapidly customize and share DMs across public platforms, it also introduces serious risks: many models are fine-tuned on unverified, proprietary, or sensitive datasets, with minimal oversight. This raises an urgent need for concept auditing: verifying whether a shared DM has learned to generate specific high-risk concepts.
@@ 3.2. Motivation and Challenges in the Wild
These challenges render the current auditing pipelines, whose reliance on observable behaviors, i.e., prompts and outputs, which is unfortunately brittle, expensive, and unreliable in real-world auditing scenarios. Therefore, we ask a more direct and scalable question: what has the model internally learned, independent of its prompts or outputs? This motivates the need for a model-centric auditing framework that avoids reliance on prompts or outputs and instead operates directly on the model’s internal behavior.
@@ 3.3.2. Theoretical Insight
The full proof is available in Appendix (Section A).
@@
Figure 2. Impact of prompts on cross-attention. We present the results of attention map $\|S\|$ and $\|diag(S)-SS^{T}\|$ with respect to the first token [BOS]. The results are averaged over all celebrity and cartoon DMs, detailed in Section 4.1.
@@
The gradient of the cross-attention function diminishes as the generation progresses (i.e., as $t$ becomes smaller).
@@ 3.3.3. Empirical Analysis
To complement this theoretical finding, we empirically examine how prompt influence evolves during the denoising process. Specifically, we track two metrics across timesteps: (1) the magnitude of the attention map $|S|$, and (2) the value of $|diag(S)-SS^{T}|$, which reflects prompt sensitivity. We focus on the [BOS] token, which typically carries the highest semantic weight in a prompt. As shown in Figure 2, in the later stages of generation (smaller $t$), the attention map becomes more focused, with $|S|$ approaching $1$. Simultaneously, $|diag(S)-SS^{T}|$ decreases toward 0, indicating a lower sensitivity to prompt variations. This empirical trend confirms our theoretical result: in the later stages of the generation process, the model becomes increasingly confident about image contents and relies less on the conditioning prompt.
@@
Note that Theorem 2 applies to all token types, showing their influence decays during denoising. While our empirical studies focus on object-level nouns for practical reasons (e.g., legal needs for auditing), the theoretical result generalizes to all tokens (nouns, adjectives, and modifiers).
@@ 3.4.1. Calibrated Error Measurement
where $\mathcal{D}_{target}$ represents the images of the target concept. Eq. 9 captures the signed difference. If the concept cannot be generated by the base model, its denoising loss is high. Fine-tuning reduces this loss, resulting in a negative $L_{ce}$ and indicating a stronger behavioral shift. Note that, given the prompt-agnostic design, the prompt $\mathbf{p}$ in Eq. 9 is not required to be accurate.
Figure 3. Calibration error (CE) across denoising time steps. We compare the difference in denoising error between the fine-tuned model and its base model for target vs. irrelevant concepts. Early stages (large $t$) offer stronger separation but are more prompt-sensitive. This motivates our conditional calibrated error (CCE), which preserves early-stage signals while reducing prompt influence.
@@
If the calibrated error falls below the threshold, the detector predicts that the fine-tuned DM can generate the target concept. Figure 3 demonstrates the effectiveness of this approach: the calibrated errors for target concepts are significantly lower than those for irrelevant concepts, enabling accurate differentiation. Note that to ensure the analysis is not influenced by inaccurate prompts, we employ the original (accurate) prompts used during fine-tuning [^4] The original prompts are used for diagnostic analysis only. All evaluations assume original prompts are unavailable.. The calibration error is normalized by the denoising error of the base model to account for variations across time steps, ensuring a fair comparison.
@@ 3.4.2. Conditional Calibrated Error Measurement
While calibrated error effectively identifies target concepts, it exhibits varying performance across different stages of the DM generation process. As shown in Figure 3, the calibration errors (using original prompts) in the early stages (large $t$) are more effective in concept identification. This finding aligns with the recent studies on DM generation mechanisms (zhang2024cross, ; yi2024towards, ): DMs primarily generate semantic information (e.g., structure) during early stages, while later stages focus on refining image details (e.g., texture). Since semantic information is more directly tied to concepts, details generated in later stages alone may lack sufficient distinction for effective concept auditing.
@@ 3.6. Overall PAIA Pipeline
We present the overall framework of PAIA, which integrates the prompt-agnostic and image-free design principles, along with unsupervised concept detector. The complete procedure of PAIA consists of the following four steps, as illustrated in Figure 4. A detailed algorithm is provided in the Appendix (Algorithm 1).
  • Step 1: Construct Irrelevant Concept Set as Baseline. Collect a set of images representing concepts the DM is not expected to generate. For each image, compute conditional calibrated error $\mathcal{L}_{cce}^{t}$ (Eq. 11) across multiple time steps. CCE integrates early-stage measurements with frozen cross-attention parameters and late-stage measurements using all fine-tuned parameters.
@@
  • Step 4: Audit Concept Generation. Apply the trained outlier detector to the target concept’s calibrated errors. If these values are statistically distinct from the baseline (i.e., flagged as outliers), the model concludes that the DM has learned to generate the target concept.
@@ 4.1. Experimental Settings
We first evaluate the performance of PAIA on two categories of fine-tuned DMs and compare it against six baselines. Details on dataset construction, image examples, and random prompt term lists are provided in Appendix Section D.
@@ 4.1.1. Datasets and Fine-tuned DMs
We generate textual prompts for each image using BLIP (li2022blip, ). Datasets are split evenly into training and testing subsets; training data is used for LoRA fine-tuning, and test data is used for auditing evaluation. The sample visualizations are included in Appendix Section D.
All DMs are fine-tuned from Stable Diffusion 1.5777https://huggingface.co/stabilityai/stable-diffusion-1-5 using Low-Rank Adaptation (LoRA), with rank $r=64$, a learning rate of 1e1e-44, and the AdamW optimizer. Each model is trained for $50$ epochs. For each category, we fine-tune models with $1$, 22, and $3$ concepts per model to evaluate multi-concept auditing. This results in $50$ celebrity and $10$ cartoon models for each setting.
@@ 4.1.2. PAIA Settings
Outlier detection model. In PAIA, we employ Isolation Forest as the detector, using the default implementation in (liu2008isolation, ) with its implicit automatic decision threshold (no additional calibration). We train the model on the irrelevant images, which cannot be generated by the DM model, and then predict if the images of the target concept are irrelevant. If not, then we predict the DM can generate this concept and vice versa.
The outlier detection model is trained with $100$ images with irrelevant concepts. In our evaluation, we randomly select images from the same category (celebrity or cartoon) but with different concepts. In practice, a straightforward strategy for constructing such a set is to use newly generated or low-popularity concepts, such as a lesser-known public figure or a recently created character.
@@
  • Random. The prompt is composed of five randomly selected terms from a common word list for generating DMs, e.g., “natural lighting,” “best quality,” “ultra detailed.” A complete list is presented in Table 10 in Appendix Section C.
  • Null. The null text, i.e., “”, is employed as the prompt input.
@@ 4.2. Baselines
  • Naive: We consider an image captioning model, GIT (wang2022git, ), to derive prompts from the target images.
@@
  • Image Classifier: We train a ConvNeXt (liu2022convnet, )-based image classifier to identify celebrity/cartoon concepts. The classifier is pre-trained on ImageNet and fine-tuned on the collected training data. After fine-tuning, the image classifier achieves 89.289.2% and 98.198.1% accuracy on celebrity and cartoon data, respectively.
@@ 4.3. Comparison with Baselines
Table 1. Performance comparison on Celebrity DMs. We report accuracy, precision, recall, F1 score, and auditing time for auditing a concept using $10$ images.
@@
We compare PAIA with baselines and show the results in Tables 1 and 2. PAIA consistently achieves the highest accuracy (92% and 91%), precision (90% and 89%), and recall (94% for both categories), demonstrating its superior performance in concept auditing. Among the baselines, the PEZ Classifier performs relatively well with accuracy of 75% and 77%, but still significantly lower than PAIA. Other methods, such as MU Classifier and MU CLIP, achieve high recall but suffer from poor precision and accuracy, indicating a high positive rate.
@@
Compared to baselines, PAIA demonstrates much greater efficiency and scalability. Prompt probing techniques (MU and PEZ) often require considerably longer optimization times. Even against naive image-captioning-based methods, PAIA demonstrates slightly higher efficiency in the second scenario since it computes denoising errors at a limited number of time steps rather than performing the full generation process.
@@ 4.4. Ablation Study
Effectiveness of Conditional Calibrated Error Measurement: We analyze the effectiveness of conditional calibrated error measurement (Eq. 11). We compare the auditing performance across four settings: 1) freezing cross-attention layers at the early stages (i.e., w/ conditional calibration), 2) not freezing any layers (i.e., w/o conditional calibration), 3) freezing cross-attention layers at all stages (i.e., w/ self-attention only), and 4) freezing self-attention layers at all stages (i.e., w/ cross-attention only). As shown in Figures 6 and 7, by freezing parameters in cross-attention layers at the early stages of generation, conditional calibration error improves the performance of PAIA in most cases.
@@
Effectiveness of Outlier Detection Algorithms: We investigate the effectiveness of different outlier detection methods, including Isolation Forest (IF) (liu2008isolation, ), Angle-based outlier detection (ABOD) (kriegel2008angle, ), k-Nearest Neighbors (kNN) (laaksonen1996classification, ), Gaussian Mixture Model (GMM) (aggarwal2017introduction, ), and One-class SVM (OneSVM) (manevitz2001one, ). We evaluate these algorithms on both celebrity and cartoon DMs with 1 concept. Among the algorithms investigated, Isolation Forest (the default algorithm in PAIA) consistently achieved the highest accuracy and F1 scores, but other robust outlier detection algorithms like ABOD and kNN also performed competitively, with slightly lower precision.
@@
Prompt Sensitivity Analysis: To quantify the prompt sensitivity in early stages, we compared auditing accuracy across three prompt strategies: caption, random, and null, where the “caption” strategy is most similar to the original prompt, compared to “random” and “null” strategies. As shown in Figure 10, auditing performance is more sensitive to prompts in early stages (large $t$), as caption prompts significantly outperform random and null prompts. In contrast, accuracy degrades at later stages (small $t$) for all prompts, where calibrated errors become less effective in auditing. These findings motivate our CCE design, which combines prompt-aware early-stage and prompt-agnostic late-stage signals.
(a) celebrity
(b) cartoon
Figure 10. Prompt sensitivity analysis in different stages.
4.5. Performance of PAIA Against Adaptive Attacks
We evaluate the robustness of PAIA under adaptive attack scenarios where attackers have full knowledge of the auditing framework and intentionally modify the fine-tuning procedure to evade detection. Specifically, we design five adaptive strategies that explicitly target the threat model and the core design of PAIA. The adaptive attacks are described below.
  • Prompt Deviation Attack: The attacker intentionally fine-tunes DMs using mismatched prompts (e.g., using “Tom Hanks” as the prompt for a model fine-tuned on “Brad Pitt” images). We train $50$ celebrity and $10$ cartoon models with deviated prompts to evaluate PAIA’s robustness.
  • Limited Calibration Attack: PAIA relies on estimating calibration errors on irrelevant data to train its outlier detector. In this attack, the auditor is restricted to a smaller irrelevant dataset, reducing the accuracy of calibration and weakening detection.
  • Regularization Attack: The attacker adds the calibrated error (Eq. 9) as a regularization term during fine-tuning. This attack aims to explicitly suppress the behavioral difference between the fine-tuned model and the base model, thereby reducing the discriminative power of PAIA and bypassing the detection.
  • Early-Freezing Attack: In this attack, the fine-tuning is restricted to the late stages (i.e., $t<\gamma$), leaving the DMs frozen in the early stages. Since early stages are more effective at revealing fine-tuning-induced behavioral shifts, by freezing the DMs during this stage, this attack aims to minimize the model changes detectable PAIA while still preserving the generation performance (by finetuning in the late stage).
  • Late-Freezing Attack: DMs are fine-tuned only during early stages ($t>\gamma$), with late (prompt-sensitive) stages frozen. By avoiding updates to the stages most influenced by specific prompts, this attack aims to keep the DMs prompt-agnostic, not limiting to the specific prompts in fine-tuning, thus reducing prompt-induced divergence between the fine-tuned and base models.
These adaptive attacks are specifically designed to exploit the key “white-box” assumption. In this setting, the auditor knows the model parameters, but not the fine-tuning process. Accordingly, the adaptive attacks could modify aspects of fine-tuning process hidden from the auditor: Prompt Deviation Attack alters the prompt in the training data; Regularization Attack modifies the training loss by regularizing the calibration error; Early-Freezing Attack and Late-Freezing Attack change which parameters are updated during fine-tuning. By targeting unseen parts of the pipeline, these attacks represent realistic adaptive strategies under practical scenarios.
Performance against Prompt Deviation Attack: As shown in Tables 5 and 6, PAIA remains unaffected by Prompt Deviation Attack, achieving a comparable performance to those with normal prompts (Tables 1 and 2). This is mainly due to our prompt-agnostic design. Across both datasets, PAIA consistently outperforms all baseline methods, achieving the highest accuracy, precision, recall, and F1 score. PEZ Classifier shows competitive performance across baselines, while methods like MU Classifier and MU CLIP continue to suffer from high false positive rates, leading to poor precision and low overall accuracy. Interestingly, these methods perform even worse than the naive prompt probing approaches, such as Naive Classifier and Naive CLIP, which do not rely on sophisticated optimization. This suggests that the MU methods face significant challenges when attempting to optimize for deviated prompts. The results underscore PAIA’s robustness against Prompt Deviation Attack.
We also observe that auditing performance under Prompt Deviation Attack is often better than under normal prompts. We attribute this to differences in the fine-tuned DMs rather than the detectors themselves, likely due to stronger memorization of deviated prompts in fine-tuned DMs, a phenomenon noted in backdoor attack research. For instance, Naive Classifier and Naive CLIP demonstrate improved auditing performance on deviated prompts, despite not utilizing any prompt information. This suggests that the difference arises from the performance of the fine-tuned DMs rather than the concept auditing methods.
Table 5. Performance comparison under Prompt Deviation Attack (Celebrity DMs). PAIA outperforms baseline detectors, while achieving comparable performance to those without adaptive attacks in Table 1.
Detector Accuracy Precision Recall F1 Score
Naive Classifier 75% 86% 60% 71%
Naive CLIP 77% 81% 70% 75%
MU Classifier 57% 54% 100% 70%
MU CLIP 57% 54% 100% 70%
PEZ Classifier 82% 81% 84% 82%
PEZ CLIP 65% 60% 92% 72%
PAIA 94% 89% 100% 94%
Table 6. Performance comparison under Prompt Deviation Attack (Cartoon DMs). PAIA outperforms baseline detectors, while achieving comparable performance to those without adaptive attacks in Table 2.
Detector Accuracy Precision Recall F1 Score
Naive Classifier 85% 86% 84% 85%
Naive CLIP 55% 55% 58% 56%
MU Classifier 56% 53% 100% 69%
MU CLIP 45% 47% 86% 61%
PEZ Classifier 83% 75% 100% 85%
PEZ CLIP 53% 52% 72% 61%
PAIA 92% 86% 100% 93%
Performance against Limited Calibration Attack: PAIA relies on accurate estimation of the conditional calibration error distributions for irrelevant images to train the outlier detector. To evaluate PAIA’s robustness against Limited Calibration Attack, we vary the number of irrelevant images used. As shown in Figure 11, $120$ images are sufficient for reliable concept auditing. This number is realistic since irrelevant examples (e.g., less popular celebrities or new cartoon characters) are typically easy for platforms to collect. Adding more images beyond this threshold yields little improvement in accuracy while increasing auditing time.
Figure 11. Robustness of PAIA against Limited Calibration Attack. We evaluate PAIA’s performance by limiting the number of irrelevant images used in the outlier detector. The default number of irrelevant images is $200$ (no attack).
Table 7. Robustness of PAIA against Regularization, Late-Freezing, and Early-Freezing Attacks. Auditing accuracy is reported on Celebrity and Cartoon DMs.
Adaptive Attacks Celebrity DMs Cartoon DMs
No Attack 92% 92%
Regularization Attack 90% 92%
Late-Freezing Attack 94% 90%
Early-Freezing Attack 93% 91%
Performance Against Regularization, Late-Freezing, and Early-Freezing Attacks: As shown in Table 7, PAIA consistently maintains high auditing accuracy across all three adaptive attacks. Specifically, the performance drops only marginally (within 2%) for all the attacks, demonstrating strong robustness of PAIA. The Regularization Attack and Late-Freezing Attack show some effectiveness in the celebrity and cartoon categories, respectively, while the Early-Freezing Attack is ineffective in both. We further analyzed the results. In the Regularization Attack, we observe that, the calibrated error (Eq. 9) is significantly reduced by the attack and becomes much smaller compared with the denoising error (Eq. 4). However, PAIA remains effective even when the behavioral difference becomes subtle. The ineffectiveness of the Early-Freezing Attack is due to PAIA’s prompt-agnostic design, which does not rely on specific training prompts and thus resists prompt-diversity obfuscation. Finally, we believe both the Early- and Late-Freezing Attacks fail to bypass PAIA due to the flexibility of the Conditional Calibrated Error, which captures behavioral shifts across both early and late stages.